OpenStudy (anonymous):

What are the length and width of a rectangle if the area is 13 1/2 square units and the length of the one side is 1/5 of measure of the perimeter.

OpenStudy (anonymous):

Perimeter of a rectangle = 2(l+b) Area of a rectangle = l*b

OpenStudy (anonymous):

Can someone explain the reasoning please.

OpenStudy (anonymous):

and according to this question, l = 1/5 of perimeter = 1/5 (2(l+b))

OpenStudy (anonymous):

I'm a little confused.

OpenStudy (anonymous):

@sauravshakya- I'm a little confused on your answer

OpenStudy (anonymous):

What is the length of the rectangle?

OpenStudy (jiteshmeghwal9):

length=\({1\over5} {\times} 2(L+B)\)

OpenStudy (jiteshmeghwal9):

now let the width be x. then \(\LARGE {(x) \times 2(L+B) \over 5}=13 \frac{1}{2}\)

OpenStudy (jiteshmeghwal9):

\(\LARGE{x \times 2L+2B \over 5}={27 \over2}\) \(\LARGE{x \times 2L + 2x={27\over2} \times 5}\)

OpenStudy (jiteshmeghwal9):

so now can u do it from here @CrazyC246 ???

OpenStudy (anonymous):

Thank you SO much @jiteshmeghwal9

OpenStudy (anonymous):

Not really.

OpenStudy (jiteshmeghwal9):

yw :) It was nice to help u

OpenStudy (jiteshmeghwal9):

well i think u should try it first again :D